OpenStudy (anonymous):

does the equation y=(0.3)^x represent exponential growth or exponential decay?

OpenStudy (anonymous):

In general for a function in the form: \(y=(a)^x\), where \(a\) is any number: If \(\{0\lt a\lt1\}\), then you have decay. If \(\{1\lt a\lt\infty\}\), then you have growth. So since \(\{0\lt0.3\lt1\}\), \(y=(0.3)^x\) represents decay

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No prob :)